OpenStudy (anonymous):

Please help with one quick question. Find the solution to the system of equations using either substitution or the linear combination method. Show all your work. 3x+5y= -9 2x+y= -34 I keep getting stuck. \=

OpenStudy (amoodarya):

multiply sec ond eq. by -5 to eliminate y and find x

OpenStudy (amoodarya):

3x+5y= -9 2x+y= -34 3x+5y= -9 -10x-5y=+170 then sum of 2 eq. -7x=170-9 -7x=161 x=-23 can you go on ?

OpenStudy (anonymous):

Then sub in -23 for x in any equation to find y, correct?

OpenStudy (anonymous):

Thank you so much for your help. :P